Last year Mr. Perkins spent 30% of his income on rent, 10% of the remainder on clothing and 3/7 of the money that remained on food. His other expenses totaled $1000 and Mr. Perkins deposited the balance, 26% of his income, in the bank. How much did Mr. Perkins earn last year?
Answer by stanbon(75887)

Last year Mr. Perkins spent 30% of his income on rent, 10% of the remainder on clothing and 3/7 of the money that remained on food. His other expenses totaled $1000 and Mr. Perkins deposited the balance, 26% of his income, in the bank. How much did Mr. Perkins earn last year?
-------
Let his income be "x".
Equation:
x - 0.30x - (3/7)(x-0.30x)-1000 = 0.26x
--------------------------------------------
0.7x - (3/7)0.7x = 0.26x + 1000
(4/7)0.7x = 0.26x + 1000
0.4x = 0.26x + 1000
0.14x = 1000
x = 1000/0.14 = $7142.86
================================
Cheers,
Stan H.